# Black-Box Access is Insufficient for Rigorous AI Audits

### v1

@@ 4. Advantages of White-Box Access
White-box algorithms produce stronger attacks. White-box algorithms allow for gradient-based optimization of adversarial inputs, which is powerful compared to simpler search methods. For example, white-box adversarial attack algorithms against vision systems typically use the gradient of the adversarial objective with respect to the input pixels to design adversarial inputs (Goodfellow et al., 2014; Papernot et al., 2016). This is much more effective for finding vulnerabilities than unguided black-box search methods. Consequently, white-box attacks are dominant in vision applications. In reinforcement learning, white-box access to a target agent also helps develop stronger adversarial attacks against it (Casper et al., 2023d; Wang et al., 2023b). For language models, optimizing adversarial inputs with gradient-based methods is more challenging because text (unlike pixels) is discrete which prevents gradient propagation. Nonetheless, there are various state-of-the-art white-box techniques for attacking language models. These include using a differentiable approximation to the process of sampling text (Wallace et al., 2019; Song et al., 2020; Guo et al., 2021), projecting adversarial embeddings onto text embeddings (Wen et al., 2023), and performing gradient-informed searches over modifications to textual changes (Ebrahimi et al., 2017; Li et al., 2018; Ren et al., 2019; Shin et al., 2020; Liu et al., 2022b; Jones et al., 2023; Zou et al., 2023a).
@@
Many black-box and grey-box attack algorithms are simply indirect or inefficient versions of white-box ones. Many black-box attacks against AI systems involve attacking a white-box model with a white-box algorithm and then testing the resulting attack on the target black-box model (Liu et al., 2016; Zhou et al., 2018). For vision models, the main motivation behind studying black-box attacks is that white-box access is not always available to attackers (Bhambri et al., 2019). Additionally, several of the most effective attacks against state-of-the-art models, such as GPT-4 and Claude-2, have simply been the result of transferring a white-box attack generated against an open-source model to the intended black-box target model (Zou et al., 2023a). Other types of black- and grey-box attack algorithms involve inefficiently estimating gradients by analyzing outputs or sampling probabilities across many queries (Grathwohl et al., 2017; Ilyas et al., 2018) when more precise gradients could be obtained trivially with white-box access.
@@
Latent space attacks help to make stronger assurances. Typically, AI systems are attacked by crafting inputs meant to make them exhibit undesirable behavior. However, input space attacks are not well-suited to diagnose certain hard-to-find issues, including high-level misconceptions (Suri et al., 2023), anomalous failures (Ziegler et al., 2022), backdoors (Chen et al., 2017; Wu et al., 2022), and deception (Christiano, 2019; Park et al., 2023a). A complementary technique for attacking systems in the input space is to relax the problem and attack their internal latent representations. The motivation of latent space attacks is that some failure modes are easier to find in the latent space than in the input space (Kumari et al., 2019; Hubinger, 2019; Jermyn, 2022; Christiano, 2019) because concepts important to the system’s reasoning are represented at a higher level of abstraction inside the model (Athiwaratkun and Kang, 2015; Johnston and Fusi, 2023; Schwinn et al., 2023; Turner et al., 2023; Zou et al., 2023b). Thus, robustness to latent attacks enables evaluators to make stronger assurances of safe worst-case performance. Latent space attacks are also more efficient to produce because they require less gradient propagation than input space attacks (Park and Lee, 2021; Qian et al., 2021), allowing for more thorough debugging work to be conducted on a limited time and computing budget. Latent space attacks are still an active area of research, but some works have emerged showing that robustness to latent space attacks effectively indicates robustness to input space attacks in vision models (Sankaranarayanan et al., 2018; Kumari et al., 2019; Zhou et al., 2019; Osada et al., 2022; Liu et al., 2023b; Zhang et al., 2023). Since textual inputs to language models are discrete, only latent space attacks allow for the direct use of gradient-based optimization, rendering them especially useful for language models (Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang and Bharti, ; Li and Qiu, 2021; Sae-Lim and Phoomvuthisarn, 2022; Pan et al., 2022; Kitada and Iyatomi, 2023).
@@
White-box methods expand the attack toolbox. Some black-box attack methods are competitive, particularly against language models. These include methods based on local search (Prasad et al., 2022), rejection sampling at scale (Ganguli et al., 2022), Langevin dynamics (Shi et al., 2022; Kumar et al., 2022), evolutionary algorithms (Lapid et al., 2023), and reinforcement learning (Deng et al., 2022; Perez et al., 2022b; Casper et al., 2023c). Additionally, some of the most effective methods for attacking language models involve human or human-guided generation of adversarial prompts (Liu et al., 2023a; Wei et al., 2023; Shah et al., 2023). However, even when black-box attacks are useful, white-box algorithms are complementary because they generate qualitatively different kinds of attacks. For example, many black-box techniques produce attacks that appear as natural language (e.g., (Shah et al., 2023)) while white-box algorithms are state-of-the-art for synthesizing nonsense adversarial prompts (e.g., (Zou et al., 2023a)), and latent space attacks produce inner perturbations. Black- and white-box methods can also be combined to conduct hybrid attacks by using the results of one method as an initialization for another. Combinations of attacks tend to be better at helping humans find vulnerabilities than a single method alone (Casper et al., 2023e).
@@
White-box interpretability tools help evaluators discover novel failure modes. White-box algorithms and interpretability tools have aided researchers in finding vulnerabilities. Examples have involved identifying internal representations of spurious features (Gandelsman et al., 2023), brittle feature representations (Carter et al., 2019; Ghorbani and Zou, 2020; Mu and Andreas, 2020; Hernandez et al., 2021; Yuksekgonul et al., 2022; Casper et al., 2022a, b, 2023e; Wu et al., 2023), and limitations of key-value memories in transformers (Geva et al., 2020; Meng et al., 2022; Geva et al., 2023). As an added benefit, attributing the problem to specific parts of the system’s architecture or representations allows developers to address it in a more precise way (Wang et al., 2023a).
@@
Consider an example. Suppose that an auditor wants to assess sycophancy: a language model’s tendency to pander to the biases of users who chat with it (Perez et al., 2022a; Sharma et al., 2023). For example, an evaluator might be concerned that the system will respond differently when the user says they are conservative or liberal in the chat. Black-box techniques could only be used to argue that the system is not sycophantic by producing examples and analyzing them for apparent sycophancy. However, a white-box interpretability-based approach could offer much more information. For example, if it were not possible for a classifier to distinguish whether the user revealed themselves to be conservative or liberal from the model’s internal representations, then this would offer stronger evidence that the system will reliably not exhibit this type of sycophancy.
Mechanistic understanding helps to make stronger assurances. In general, it is impossible to make guarantees about black-box systems using a finite number of queries without additional assumptions. In contrast to black-box methods, which can only show the existence of failures by finding inputs that elicit them, thoroughly understanding the computations inside of a model gives auditors a complementary way to find evidence against the existence of failure modes. A mechanistic understanding can help researchers develop a predictive model of how the system would act for broad classes of inputs. Some works have aimed to provide thorough investigations of how networks perform simple tasks (Cammarata et al., 2020; Nanda et al., 2023; Zhong et al., 2023). Although scaling thorough analysis is an open challenge (Tegmark and Omohundro, 2023), it offers a strategy for making strong assurances. Recent works have attempted to make progress on this problem by using sparse autoencoders to allow evaluators to more thoroughly study the features represented inside of large language models (Cunningham et al., 2023; Bricken et al., 2023).
@@
State-of-the-art AI systems are typically trained on large amounts of internet data, often in multiple stages. This can cause them to learn undesirable capabilities, such as knowledge of how to perform illegal activities (Shah et al., 2023; Zou et al., 2023b) or the ability to produce harmful content (Geirhos et al., 2020b; Birhane et al., 2021; Rando et al., 2022; Birhane et al., 2023; Thiel, 2023; Thiel et al., 2023; Shahbazi et al., 2023). Developers attempt to remove harmful abilities through fine-tuning, but they can unexpectedly resurface through “jailbreaks” (Albert, 2023; Oneal, 2023; Li et al., 2023; Liu et al., 2023a; Rao et al., 2023; Wei et al., 2023; Shen et al., 2023; Qi et al., 2023b; Yu et al., 2023; Deng et al., 2023b; Yong et al., 2023; Zou et al., 2023a; Chao et al., 2023; Shah et al., 2023; Tsai et al., 2023) or further fine-tuning models on a small number of new examples (Yang et al., 2023; Qi et al., 2023a; Lermen et al., 2023; Zhan et al., 2023). The existence of harmful dormant capabilities in models thus poses risks from attacks and fine-tuning, especially if they are leaked (e.g. Stable Diffusion (Lanz, 2023)), open-sourced (e.g., Llama-2 (Touvron et al., 2023)), or deployed with fine-tuning access via API (e.g., GPT-3.5 (OpenAI, 2023b)). Consequently, being able to fine-tune the model offers another strategy to search for evidence of undesirable capabilities and assess the risks in deployment.
@@ 5. Advantages of Outside-the-Box Access
Access to methodological details helps to identify risks. Knowing methodological details can reveal shortcuts taken during development, which can guide evaluators toward discovering problems. For example, if a system was trained with human-generated data using a non-representative cohort of humans, this can suggest specific social biases that the system may have internalized (Santurkar et al., 2023; Sun et al., 2023; Casper et al., 2023b). Knowing the findings of internal evaluations is especially useful for helping auditors target their efforts toward a set of complementary evaluations. Furthermore, when developers attempt to mitigate flaws, auditors can better assess the effectiveness of these efforts if they have detailed information about the attempted mitigation (e.g., fine-tuning datasets, both old and new versions of model weights, etc.) (Pozzobon et al., 2023).
@@
Access to data helps auditors trace problems and assess fair use. Recent work has highlighted the ability of dataset audits to identify harmful and biased content used to train models (Geirhos et al., 2020b; Birhane et al., 2021, 2023; Thiel, 2023; Thiel et al., 2023; Shahbazi et al., 2023). Access to training data also helps to investigate risks of data-poisoning attacks (which is especially important for systems trained on internet data) (Carlini et al., 2023b; Rando and Tramèr, 2023; Wang et al., 2023c; Wan et al., 2023). Meanwhile, legal questions are currently being debated involving the extent to which training generative AI systems on copyrighted content constitutes fair use (Karamolegkou et al., 2023; Henderson et al., 2023; Rodriguez Maffioli, 2023; Company, 2023). Auditors may require access to training data to properly assess whether it was used in accordance with copyright law.
@@
Contextual information makes it easier to hold developers accountable. Requirements to produce documentation place greater responsibility on developers to detail their methods, especially if subject to regulatory penalty defaults (Wieringa, 2020; Yew and Hadfield-Menell, 2022). Contextual information provides information about whether developers made decisions in a responsible manner (Mökander and Floridi, 2021). For example, documentation can provide insights into why certain design choices were made over others (Metcalf et al., 2022). Datasets and training details can help trace risks to intentional choices, and internal evaluation reports provide insights into how the developer responded to findings. By increasing the scrutiny placed on decisions in the development process, requirements for greater methodological transparency to auditors can deter developers from taking risks in the first place (Casper et al., 2023b; Lambert et al., 2023).
6. Security concerns are nonunique and addressable.
@@
A concern with white- and outside-the-box audits is an increased risk that a developer’s models or intellectual property could be leaked (Bommasani et al., 2023). In turn, leaks could compromise developers’ trade secrets and pose risks to the public if they enable misuse (Nevo et al., 2023). Widespread norms for secure audits in AI do not yet exist as there are in other industries. However, the risk of leaks can be minimized through several technical, physical, and legal mechanisms. With these measures, developers can provide white- and outside-the-box access to auditors without the system’s parameters leaving their servers. These can reduce leakage risks to a level comparable to ones posed by common existing practices.
@@
Technical: API access can offer remote auditors de facto white-box access. Forms of structured access, particularly research application programming interfaces (APIs) (Bucknall and Trager, 2023; Shevlane, 2022; Bluemke et al., 2023; Fiotto-Kaufmann et al., 2023), could enable auditors to analyze systems using some white-box tools without giving auditors direct access to model parameters. We refer to this form of access as de facto white-box access if it enables auditors to indirectly run arbitrary white-box processes on models while restricting direct access to model parameters. One example of running an algorithm that accesses a model’s parameters via API is an OpenAI GPT-3.5 API which allows for fine-tuning (OpenAI, 2023b). However, more customizable APIs (e.g., (Fiotto-Kaufmann et al., 2023)) would be needed to allow for more flexible access. Another proposed paradigm is a flexible query API (Openmined, 2023), where auditors are given complete access to mock versions of a model and data. The auditors then develop evaluations using their complete access to these mock artifacts before submitting them to be run on the true model and dataset. This allows auditors to better customize their evaluations.
@@
Physical: Secure research environments can be used for auditors given unrestricted white-box access. Auditing personnel could securely be given white-box access to a system by hosting them on-site at the developer’s facilities in a secure research environment. This is a common practice in other industries despite the costs of requiring auditors to be physically on-site (van den Brom, 2022; Openmined, 2023). Compared to API access, this could allow auditors to access systems more flexibly and efficiently while minimizing the risk that the model is leaked or reconstructed. Safeguards for limiting information leakage through lab employees (such as NDAs) are already common practice and could be adapted for application to on-site auditors. However, it is unclear whether, absent external legal structures, labs could incentivize adherence to protective measures to the same extent as with employees.
Legal: Auditors in other industries have developed practices to address the risks of leaks. Across many industries, auditors require privileged access to systems and data in order to perform effective assessments. There are established mechanisms from other fields, such as financial auditing, employed to reduce the risk of leaks. In the finance industry, this manifests in three main ways. First, policies for confidentiality and handling of sensitive information are enforced by formal training and non-disclosure clauses in contracts to hold auditors accountable for violations (EY, 2019). Second, there are clear terms of engagement that govern the relationship between auditor and auditee. These typically include specific restrictions on confidentiality expectations tailored to a particular client (ASA, 2006). While some specifics of auditing are managed through contracting, the Public Company Accounting Oversight Board (PCAOB) requires all registered auditors to adhere to common standards in the US (PCAOB, 2002). Finally, auditors can be legally required to avoid conflicts of interest. In the US, this is done through a regime specifying general provisions for auditor independence (such as reporting requirements) outlined in Title II of the Sarbanes Oxley Act (PCAOB, 2002). Provisions are enforced through various agencies, including the Securities Exchange Commission (SEC), which prevents auditor manipulation or the use of financial information for personal gain (of Federal Regulations, 2023). This type of enforcement could allow for auditors to be held accountable in a way that could reduce the risk of leakage to a level comparable to risks posed by the developer’s employees. In fact, employees may pose greater risks of sharing tacit knowledge with competitors than auditors because developers regularly attempt to recruit AI researchers from competitors.

### v2

@@ 4. Advantages of White-Box Access
White-box algorithms produce stronger attacks. White-box algorithms allow for gradient-based optimization of adversarial inputs, which is powerful compared to simpler search methods. For example, white-box adversarial attack algorithms against vision systems typically use the gradient of the adversarial objective with respect to the input pixels to design adversarial inputs (Goodfellow et al., 2014; Papernot et al., 2016). This is much more effective for finding vulnerabilities than unguided black-box search methods. Consequently, white-box attacks are dominant in vision applications. In reinforcement learning, white-box access to a target agent also helps develop stronger adversarial attacks against it (Casper et al., 2023c; Wang et al., 2023a). For language models, optimizing adversarial inputs with gradient-based methods is more challenging because text (unlike pixels) is discrete which prevents gradient propagation. Nonetheless, there are various state-of-the-art white-box techniques for attacking language models. These include using a differentiable approximation to the process of sampling text (Wallace et al., 2019; Song et al., 2020; Guo et al., 2021), projecting adversarial embeddings onto text embeddings (Wen et al., 2023), and performing gradient-informed searches over modifications to textual changes (Ebrahimi et al., 2017; Li et al., 2018; Ren et al., 2019; Shin et al., 2020; Liu et al., 2022b; Jones et al., 2023; Zou et al., 2023b).
@@
Many black-box and grey-box attack algorithms are simply indirect or inefficient versions of white-box ones. Many black-box attacks against AI systems involve attacking a white-box model with a white-box algorithm and then testing the resulting attack on the target black-box model (Liu et al., 2016; Zhou et al., 2018). For vision models, the main motivation behind studying black-box attacks is that white-box access is not always available to attackers (Bhambri et al., 2019). Additionally, several of the most effective attacks against state-of-the-art models, such as GPT-4 and Claude-2, have simply been the result of transferring a white-box attack generated against an open-source model to the intended black-box target model (Zou et al., 2023b). Other types of black- and grey-box attack algorithms involve inefficiently estimating gradients by analyzing outputs or sampling probabilities across many queries (Grathwohl et al., 2017; Ilyas et al., 2018) when more precise gradients could be obtained trivially with white-box access.
@@
Latent space attacks help to make stronger assurances. Typically, AI systems are attacked by crafting inputs meant to make them exhibit undesirable behavior. However, input space attacks are not well-suited to diagnose certain hard-to-find issues, including high-level misconceptions (Suri et al., 2023), anomalous failures (Ziegler et al., 2022), backdoors (Chen et al., 2017; Wu et al., 2022), and deception (Christiano, 2019; Park et al., 2023a). A complementary technique for attacking systems in the input space is to relax the problem and attack their internal latent representations. The motivation of latent space attacks is that some failure modes are easier to find in the latent space than in the input space (Casper et al., 2024) because concepts important to the system’s reasoning are represented at a higher level of abstraction inside the model (Athiwaratkun and Kang, 2015; Johnston and Fusi, 2023; Schwinn et al., 2023; Turner et al., 2023; Zou et al., 2023a). Thus, robustness to latent attacks enables evaluators to make stronger assurances of safe worst-case performance. Latent space attacks are also more efficient to produce because they require less gradient propagation than input space attacks (Park and Lee, 2021; Qian et al., 2021), allowing for more thorough debugging work to be conducted on a limited time and computing budget. Latent space attacks are still an active area of research, but some works have emerged showing that robustness to latent space attacks effectively indicates robustness to input space attacks in vision models (Sankaranarayanan et al., 2018; Kumari et al., 2019; Zhou et al., 2019; Osada et al., 2022; Liu et al., 2023b; Zhang et al., 2023). Since textual inputs to language models are discrete, only latent space attacks allow for the direct use of gradient-based optimization, rendering them especially useful for language models (Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang and Bharti, [n. d.]; Li and Qiu, 2021; Sae-Lim and Phoomvuthisarn, 2022; Pan et al., 2022; Kitada and Iyatomi, 2023).
@@
White-box methods expand the attack toolbox. Some black-box attack methods are competitive, particularly against language models. These include methods based on local search (Prasad et al., 2022), rejection sampling at scale (Ganguli et al., 2022), Langevin dynamics (Shi et al., 2022; Kumar et al., 2022), evolutionary algorithms (Lapid et al., 2023), and reinforcement learning (Deng et al., 2022; Perez et al., 2022a; Casper et al., 2023e). Additionally, some of the most effective methods for attacking language models involve human or human-guided generation of adversarial prompts (Liu et al., 2023a; Wei et al., 2023; Shah et al., 2023). However, even when black-box attacks are useful, white-box algorithms are complementary because they generate qualitatively different kinds of attacks. For example, many black-box techniques produce attacks that appear as natural language (e.g., (Shah et al., 2023)) while white-box algorithms are state-of-the-art for synthesizing adversarial prompts which appear to humans as unintelligible text (e.g., (Zou et al., 2023b)). Relying only on black-box methods would likely miss out on these types of failures entirely because finding them with black-box searches or human inductive biases would be highly improbable. Black- and white-box methods can also be combined to conduct hybrid attacks by using the results of one method as an initialization for another. Combinations of attacks tend to be better at helping humans find vulnerabilities than a single method alone (Casper et al., 2023d).
@@
White-box interpretability tools help evaluators discover novel failure modes. White-box algorithms and interpretability tools have aided researchers in finding vulnerabilities. Examples have involved identifying novel attacks (Ziegler et al., 2022), internal representations of spurious features (Gandelsman et al., 2023; Marks et al., 2024), brittle feature representations (Carter et al., 2019; Ghorbani and Zou, 2020; Mu and Andreas, 2020; Hernandez et al., 2021; Yuksekgonul et al., 2022; Casper et al., 2022b, a, 2023d; Wu et al., 2023), and limitations of key-value memories in transformers (Geva et al., 2020; Meng et al., 2022; Geva et al., 2023). As an added benefit, attributing the problem to specific parts of the system’s architecture or representations allows developers to address it in a more precise way (Wang et al., 2023c).
@@
Consider an example. Suppose that an auditor wants to assess sycophancy: a language model’s tendency to pander to the biases of users who chat with it (Perez et al., 2022b; Sharma et al., 2023). For example, an evaluator might be concerned that the system will respond differently when the user says they are conservative or liberal in the chat. Black-box techniques could only be used to argue that the system is not sycophantic by producing examples and analyzing them for apparent sycophancy. However, a white-box interpretability-based approach could offer much more information. For example, if it were not possible for a classifier to distinguish whether the user revealed themselves to be conservative or liberal from the model’s internal representations, then this would offer stronger evidence that the system will reliably not exhibit this type of sycophancy.
Mechanistic understanding helps to make stronger assurances. In general, it is impossible to make guarantees about black-box systems using a finite number of queries without additional assumptions. In contrast to black-box methods, which can only show the existence of failures by finding inputs that elicit them, thoroughly understanding the computations inside of a model gives auditors a complementary way to find evidence against the existence of failure modes. A mechanistic understanding can help researchers develop a predictive model of how the system would act for broad classes of inputs. Some works have aimed to provide thorough investigations of how networks perform simple tasks (Cammarata et al., 2020; Nanda et al., 2023; Zhong et al., 2023). Although scaling thorough analysis is an open challenge (Tegmark and Omohundro, 2023), it offers a strategy for making strong assurances. Recent works have attempted to make progress on this problem by using sparse autoencoders to allow evaluators to more thoroughly study the features represented inside of large language models (Cunningham et al., 2023; Bricken et al., 2023; Marks et al., 2024).
@@
State-of-the-art AI systems are typically trained on large amounts of internet data, often in multiple stages. This can cause them to learn undesirable capabilities, such as knowledge of how to perform illegal activities (Shah et al., 2023; Zou et al., 2023a) or the ability to produce harmful content (Geirhos et al., 2020b; Luccioni and Viviano, 2021; Birhane et al., 2021; Rando et al., 2022; Birhane et al., 2023; Thiel, 2023; Thiel et al., 2023; Shahbazi et al., 2023). Developers attempt to remove harmful abilities through fine-tuning, but they can unexpectedly resurface through “jailbreaks” (Albert, 2023; Oneal, 2023; Li et al., 2023; Liu et al., 2023a; Rao et al., 2023; Wei et al., 2023; Shen et al., 2023; Qi et al., 2023a; Yu et al., 2023; Deng et al., 2023a; Yong et al., 2023; Zou et al., 2023b; Chao et al., 2023; Shah et al., 2023; Tsai et al., 2023) or further fine-tuning models on a small number of new examples (Yang et al., 2023; Qi et al., 2023b; Lermen et al., 2023; Zhan et al., 2023). The existence of harmful dormant capabilities in models thus poses risks from attacks and fine-tuning, especially if they are leaked (e.g. Stable Diffusion (Lanz, 2023)), open-sourced (e.g., Llama-2 (Touvron et al., 2023)), or deployed with fine-tuning access via API (e.g., GPT-3.5 (OpenAI, 2023a)). Consequently, being able to fine-tune the model offers another strategy to search for evidence of undesirable capabilities and assess the risks in deployment.
@@ 5. Advantages of Outside-the-Box Access
Access to methodological details helps to identify risks. Knowing methodological details can reveal shortcuts taken during development, which can guide evaluators toward discovering problems. For example, if a system was trained with human-generated data using a non-representative cohort of humans, this can suggest specific social biases that the system may have internalized (Santurkar et al., 2023; Sun et al., 2023; Casper et al., 2023a). Knowing the findings of internal evaluations is especially useful for helping auditors target their efforts toward a set of complementary evaluations. Furthermore, when developers attempt to mitigate flaws, auditors can better assess the effectiveness of these efforts if they have detailed information about the attempted mitigation (e.g., fine-tuning datasets, both old and new versions of model weights, etc.) (Pozzobon et al., 2023).
@@
Access to data helps auditors trace problems and assess fair use. Recent work has highlighted the ability of dataset audits to identify harmful and biased content used to train models (Geirhos et al., 2020b; Luccioni and Viviano, 2021; Birhane et al., 2021, 2023; Thiel, 2023; Thiel et al., 2023; Shahbazi et al., 2023). For example, Birhane et al. (2021) and Thiel (2023) were able to identify previously-overlooked examples of unintended hateful, sexual, and child-abuse-related content in widely-used training datasets. Access to training data also helps to investigate risks of data-poisoning attacks (which is especially important for systems trained on internet data) (Carlini et al., 2023a; Rando and Tramèr, 2023; Wang et al., 2023b; Wan et al., 2023). Meanwhile, legal questions are currently being debated involving the extent to which training generative AI systems on copyrighted content constitutes fair use (Karamolegkou et al., 2023; Henderson et al., 2023; Rodriguez Maffioli, 2023; Company, 2023). Auditors may require access to training data to properly assess whether it was used in accordance with copyright law.
@@
Contextual information makes it easier to hold developers accountable. Requirements to produce documentation place greater responsibility on developers to detail their methods, especially if subject to regulatory penalty defaults (Wieringa, 2020; Yew and Hadfield-Menell, 2022). Contextual information provides information about whether developers made decisions in a responsible manner (Mökander and Floridi, 2021). For example, documentation can provide insights into why certain design choices were made over others (Metcalf et al., 2022). Datasets and training details can help trace risks to intentional choices, and internal evaluation reports provide insights into how the developer responded to findings. By increasing the scrutiny placed on decisions in the development process, requirements for greater methodological transparency to auditors can deter developers from taking risks in the first place (Casper et al., 2023a; Lambert et al., 2023).
6. Methods to Address Security Risks
@@
A concern with white- and outside-the-box audits is an increased risk that a developer’s models or intellectual property could be leaked (Bommasani et al., 2023). In turn, leaks could compromise developers’ trade secrets and pose risks to the public if they enable misuse (Nevo et al., 2023). Widespread norms for secure audits in AI do not yet exist, but there is precedent in other fields for navigating similar challenges to enable secure oversight. The risk of leaks can be minimized through several technical, physical, and legal mechanisms. With these measures, developers can provide white- and outside-the-box access to auditors without the system’s parameters leaving their servers. These can reduce leakage risks to a level comparable to ones posed by common existing practices.
@@
Technical: API access can offer remote auditors de facto white-box access. Forms of structured access, particularly research application programming interfaces (APIs) (Bucknall and Trager, 2023; Shevlane, 2022; Bluemke et al., 2023; Fiotto-Kaufmann et al., 2023), could enable auditors to analyze systems using some white-box tools without giving auditors direct access to model parameters. We refer to this form of access as de facto white-box access if it enables auditors to indirectly run arbitrary white-box processes on models while restricting direct access to model parameters. One example of running an algorithm that accesses a model’s parameters via API is an OpenAI GPT-3.5 API which allows for fine-tuning (OpenAI, 2023a). However, more customizable APIs (e.g., (Fiotto-Kaufmann et al., 2023)) would be needed to allow for more flexible access. Another proposed paradigm is a flexible query API (Openmined, 2023), where auditors are given complete access to mock versions of a model and data. The auditors then develop evaluations using their complete access to these mock artifacts before submitting them to be run on the true model and dataset. This allows auditors to better customize their evaluations.
@@
Physical: Secure research environments can be used for auditors given unrestricted white-box access. Auditing personnel could securely be given white-box access to a system by hosting them on-site at the developer’s facilities in a secure research environment. This is a common practice in other industries despite the costs of requiring auditors to be physically on-site (van den Brom, 2022; Openmined, 2023; Trager et al., 2023). For example, the International Atomic Energy Agency employs over 300 expert inspectors from approximately 80 countries who do on-site inspections of nuclear facilities (International Atomic Energy Agency, 2016, 2023). Compared to API access, secure research environments could allow auditors to access systems more flexibly and efficiently while minimizing the risk that the model is leaked or reconstructed. Safeguards for limiting information leakage through lab employees (such as NDAs) are already common practice and could be adapted for application to on-site auditors. However, it is unclear whether, absent external legal structures, labs could incentivize adherence to protective measures to the same extent as with employees.
Legal: Other industries have developed practices to address the risk of leaks from audits. Across many industries, auditors require privileged access to systems and data in order to perform effective assessments. There are established mechanisms from other fields, such as financial auditing, employed to reduce the risk of leaks. In the finance industry, this manifests in three main ways. First, policies for confidentiality and handling of sensitive information are enforced by formal training and non-disclosure clauses in contracts to hold auditors accountable for violations (EY, 2019). Second, there are clear terms of engagement that govern the relationship between auditor and auditee. These typically include specific restrictions on confidentiality expectations tailored to a particular client (ASA, 2006). While some specifics of auditing are managed through contracting, the Public Company Accounting Oversight Board (PCAOB) requires all registered auditors to adhere to common standards in the US (PCAOB, 2002). Finally, auditors can be legally required to avoid conflicts of interest. In the US, this is done through a regime specifying general provisions for auditor independence (such as reporting requirements) outlined in Title II of the Sarbanes Oxley Act (PCAOB, 2002). Provisions are enforced through various agencies, including the Securities Exchange Commission (SEC), which prevents auditor manipulation or the use of financial information for personal gain (of Federal Regulations, 2023). This type of enforcement could allow for auditors to be held accountable in a way that could reduce the risk of leakage to a level comparable to risks posed by the developer’s employees. In fact, employees may pose greater risks of sharing tacit knowledge with competitors than auditors because developers regularly attempt to recruit AI researchers from competitors.